Problem 6.23 (Solution Video) Your answer is incorrect. Try again. Sunland Energy Company owns several gas stations. Management is looking to open a new station in the western suburbs of Baltimore. One possibility that managers at the company are evaluating is to take over a station located at a site that has been leased from the county. The lease, originally for 99 years, currently has 73 years before expiration. The gas station generated a net cash flow of $93,580 last year, and the current owners expect an annual growth rate of 6.3 percent. If Sunland Energy uses a discount rate of 13.6 percent to evaluate such businesses, what is the present value of this growing annuity? (Round factor values to 6 decimal places, e.g. 1.521253 and final answer to 2 decimal places, e.g. 15.21.) aceound factor Present val,271,867. Question Attempts: 1 of 2 used SAVE FOR LATER SUBMIT ANSWE
Answer #1

The PV of Growing Annuity = Cash Flow*(1+growth)/(1+rate) + Cash Flow*((1+growth)/(1+rate))2 + Cash Flow*((1+growth)/(1+rate))3 + ......... Cash Flow*((1+growth)/(1+rate))73
PV of Growing Annuity = 93,850*(1.063/1.136) + 93,850*(1.063/1.136)2 + 93,850*(1.063/1.136)3 ........93,850*(1.063/1.136)73
= 92850 *1.063/1.136 * (1- (1.063/1.136)73)/(1-(1.063/1.136)) = 1,355,896.196306

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well